OpenStudy (jamaljunior):

Given the equation P = 2w + 2l solve for w.

jhonyy9 (jhonyy9):

like a first step subtract from both sides 2l so what will get ?

OpenStudy (vuriffy):

P = 2w + 2l You must get w by itself. Subtract 2l P - 2l = 2w What do you think you should do next?

OpenStudy (jamaljunior):

what is p?

OpenStudy (vuriffy):

We are just dealing with variables, it won't do anything really.

OpenStudy (vuriffy):

It's like if we were looking for what P equaled, that would be.

jhonyy9 (jhonyy9):

Jamal there is given that P = 2w -2l so this is P

OpenStudy (vuriffy):

How many w's does it say you have?

OpenStudy (vuriffy):

P - 2l = 2w

OpenStudy (jamaljunior):

2

OpenStudy (vuriffy):

So you can divide by 2 to isolate w.

OpenStudy (vuriffy):

\[w = \frac{ P - 2l }{ 2 }\]

OpenStudy (jamaljunior):

idk how to divide by 2 if i dont know what P is??

OpenStudy (vuriffy):

It's just a variable.

OpenStudy (vuriffy):

You don't need to know what they actually are, numerical wise.

OpenStudy (jamaljunior):

so how to divide

OpenStudy (vuriffy):

Just by 2 to isolate w. There's 2w, so you divide by 2, and w stays isolated.

OpenStudy (vuriffy):

Look up a little and you will see what w equals and that's your answer.

OpenStudy (jamaljunior):

ohh ty